Using logrithms as required to perform the following operations.

(274,000) 2/5

Answer by stanbon(75887)

Let x = (274,000)^ (2/5)
Then log x = (2/5)log274500
log x = (2/5)5.5385...=2.1754...
Take the anti-log to get:
x = 149.767279....
==========
Cheers,
Stan H.